Two planes which are 2160 miles apart fly toward each other. Their speeds differ by 70 mph. If they pass each other in 3 hours, what is the speed of each?
Answer by mananth(16949)

Plane A speed = x mph
Plane B speed = x+70
They are travelling towards each other
combined speed = x+x+70= 2x+70
Time = 3 hours
distance = 2160
time = d/r
3=2160/(2x+70)
3(2x+70)=2160
6x+210=2160
6x=1950
x= 1950/6
x= 325 mph speed of one plane
speed of other plane = 325+70 = 395 mph
m.ananth@hotmail.ca